OpenStudy (anonymous):

Solve: 21x - 24x2 + 6x = 9 -24x2 1/2 1/3 3/2 2/3

OpenStudy (anonymous):

21x - 24x2 + 6x = 9 -24x2 21x -24x^2 + 24x^2 + 5x = 9 21x + 5x = 9 try solving from here.

OpenStudy (anonymous):

Would it be 1/3?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

That doesn't make sense to me.

OpenStudy (anonymous):

add the 21 and the 5 21x + 5x = 9 26x = 9 divide both sides by 26 to get x by itself.

OpenStudy (anonymous):

I still don't get it why wouldn't it be 1/3?